# Plug-and-Play Fidelity Optimization for Diffusion Transformer Acceleration via Cumulative Error Minimization

### v1

@@ Abstract
Although Diffusion Transformer (DiT) has emerged as a predominant architecture for image and video generation, its iterative denoising process results in slow inference, which hinders broader applicability and development. Caching-based methods achieve training-free acceleration, while suffering from considerable computational error. Existing methods typically incorporate error correction strategies such as pruning or prediction to mitigate it. However, their fixed caching strategy fails to adapt to the complex error variations during denoising, which limits the full potential of error correction. To tackle this challenge, we propose a novel fidelity-optimization plugin for existing error correction methods via cumulative error minimization, named CEM. CEM predefines the error to characterize the sensitivity of model to acceleration jointly influenced by timesteps and cache intervals. Guided by this prior, we formulate a dynamic programming algorithm with cumulative error approximation for strategy optimization, which achieves the caching error minimization, resulting in a substantial improvement in generation fidelity. CEM is model-agnostic and exhibits strong generalization, which is adaptable to arbitrary acceleration budgets. It can be seamlessly integrated into existing error correction frameworks and quantized models without introducing any additional computational overhead. Extensive experiments conducted on nine generation models and quantized methods across three tasks demonstrate that CEM significantly improves generation fidelity of existing acceleration models, and outperforms the original generation performance on FLUX.1-dev, PixArt-$\alpha$, StableDiffusion1.5 and Hunyuan. The code will be made publicly available.
@@ 1 Introduction
Diffusion models ho2020denoising; rombach2022high have significantly advanced visual generation tasks, including image chen2023pixart and video zheng2024open generation, and even extended to text generation nie2025large; hu2025accelerating. Diffusion Transformers (DiT) peebles2023scalable have emerged as the dominant architecture for diffusion models, replacing U-Net ronneberger2015u due to their inherent scalability and superior generative capabilities. Despite the impressive performance of these powerful models, their slow inference speed remains a critical barrier to widespread adoption. Currently, image generation typically requires tens of seconds, while video generation can take up to several minutes or even longer. This limitation primarily stems from the sequential nature of the reverse denoising process, which precludes parallel decoding. Moreover, the expansion of parameters, coupled with computationally intensive operations like attention vaswani2017attention, further worsens efficiency.
To accelerate the DiT-based diffusion generation, researchers have adopted general techniques such as distillation yin2024one; salimans2022progressive; gu2025video, learning huang2024harmonica; ma2024learning and quantization shang2023post; he2023ptqd; li2025dvd, aiming to reduce inference latency by decreasing the model size. However, they necessitate training phases, which entails substantial computational cost, and lack generalization across models. An alternative technical approach is the caching mechanism liu2025timestep; saghatchian2025cached; qiu2025accelerating, which leverages the similarity ma2024deepcache between adjacent timesteps or layers to reuse previous cached hidden states, achieving training-free acceleration. Naive caching selvaraju2024fora; ma2024deepcache inevitably accumulates noise throughout the denoising process, with error growing exponentially as the cache interval increases, resulting in a substantial deterioration in generation fidelity.
Therefore, existing caching optimization methods are combined with additional mechanisms to perform error correction, aiming to mitigate the cumulative caching error. For example, methods such as ToCa zou2024accelerating, DuCa zou2024accelerating1, and FastCache liu2025fastcache incorporate the pruning operation when reusing cache to retain the computation of a subset of important tokens, thereby reducing error. In addition, methods such as ICC chen2025accelerating and TaylorSeer liu2025reusing; guan2025forecasting leverage historical trends to predict the output variations across different timesteps and guide the update of cached representations, rather than directly reusing cache.
@@
Although the idea of error correction partially reduces the cumulative error for balancing the generation quality and acceleration efficiency, its effectiveness is limited by the error of relatively simple or fixed caching strategies. For example, ToCa zou2024accelerating and DuCa zou2024accelerating1 adopt the cache schedule that changes linearly with the timestep, while the cache interval of methods like ICC chen2025accelerating and TaylorSeer liu2025reusing; guan2025forecasting is a constant. They are incapable of handling the complex dynamics of sensitivity to caching during denoising. This limitation results in an insufficient characterization of the model’s intrinsic sensitivity to caching, thereby preventing the mitigation of error accumulation during acceleration. Consequently, the potential of error correction is limited, leading to degradation in generation fidelity.
@@
Although several prior works qiu2025accelerating like TeaCache liu2025timestep have explored caching optimization, they rely on real-time error estimation, introducing additional computational overhead that undermines efficiency gains. Furthermore, they are tied to specific acceleration ratios or model architectures, limiting compatibility with error correction. In contrast, our CEM conducts offline error modeling prior to inference, enabling the use of prior knowledge without incurring runtime cost. CEM is model-agnostic, supports arbitrary acceleration budgets, and integrates seamlessly with both error correction methods and quantized models.
@@
  • We introduce dynamic programming based on prior errors to derive the optimal caching strategy that minimizes cumulative error, thereby substantially reducing the error of existing caching-based methods.
@@ 2 Related Work
Diffusion Transformer. Diffusion models ho2020denoising; rombach2022high have demonstrated remarkable capabilities in image chen2023pixart and video zheng2024open; kong2024hunyuanvideo generation. Upon this, DiT peebles2023scalable replace the convolutional modules in U-Net ronneberger2015u with attention vaswani2017attention, which enhances scalability and leads to breakthrough improvements in generation quality. Despite these advancements, the iterative nature of the diffusion combined with the computational complexity of attention results in slow inference. As a result, acceleration techniques have emerged as a key focus of research.
@@
DiT Acceleration. General acceleration techniques such as distillation meng2023distillation; yao2024fasterdit and quantization li2023q; liu2025cachequant typically require retraining the models, which incur substantial time costs and lacks generalization across different generation models. Consequently, training-free methods have emerged as a promising direction. Caching-based methods selvaraju2024fora; zou2025exposure; lv2024fastercache accelerate inference by reusing previous hidden states with high similarity, while the reuse of cache inevitably introduces error, which accumulate over time and result in a significant degradation of generation quality. To correct the caching error, two strategies have been proposed: (1) Token pruning, which retains the computation of a subset of important tokens during cache reuse to reduce the error caused by fully relying on cache. For example, DuCa zou2024accelerating1 applies pruning at specific timesteps when reusing the cache to balance generation quality and acceleration efficiency. (2) Predictive reuse, which estimates future representations based on the historical evolution of cached features rather than directly reusing them. For instance, TaylorSeer liu2025reusing; guan2025forecasting performs a Taylor expansion on intermediate outputs and utilizes gradient dynamics to predict features for reuse.
@@
Although these methods alleviate caching error to some extent, they overlook the optimization of the caching strategy itself. Consequently, corrections are applied on top of relatively high cumulative error, limiting their effectiveness. Although a few studies have explored caching strategy optimization, they are difficult to integrate with error correction for generation quality enhancement. For instance, AdaCache kahatapitiya2024adaptive relies on motion trajectories and is restricted to video; AdaptiveDiffusion ye2024training introduces overhead from third-order difference calculations and is specific to the U-Net; TeaCache liu2025timestep performs real-time input-output estimation and polynomial fitting, while the extra cost offsets the benefits of cache-based acceleration. To address these issues, we propose a plug-and-play acceleration framework that optimizes the caching strategy by offline prior, seamlessly integrates with most error correction methods, and introduces no additional computational overhead, substantially improving their generation fidelity.
@@
In addition, optimizing the sampling strategy in diffusion also leads to acceleration. For example, DDIM song2020denoising introduces deterministic sampling to reduce the number of denoising steps while maintaining generation quality. DPM-Solver lu2022dpm and DPM-Solver++ lu2025dpm leverage adaptive high-order solvers to accelerate denoising. Rectified flow liu2022flow; liu2022rectified improves the transport of the distribution in the ordinary differential equation. Our method is orthogonal to these approaches and can be integrated and generalized across different models and sampling strategies, even supporting compatibility with quantized models.
@@ 3.1 Offline Error Modeling
Joint modeling analysis. We model the error as a joint function of denoising timesteps and cache intervals. Previous methods zou2024accelerating; zou2024accelerating1; liu2025reusing only consider differences across timesteps and determine cache intervals heuristically, limiting control over the acceleration budget. In contrast, we explicitly model cache intervals together with timesteps to capture the error distribution and analyze the sensitivity of different denoising stages. This formulation is simple yet accurate, offering flexible acceleration control, while its offline design introduces no additional computational overhead and provides the foundation for our algorithm of cumulative error minimization in Sec. 3.2.
@@ 4.1 Experiment Settings
Models and baselines. We conduct extensive experiments across three tasks: text-to-image generation (using StableDiffusion1.5 (SD1.5) rombach2022high, PixArt-$\bm{\alpha}$ chen2023pixart and FLUX.1-dev flux2024), text-to-video generation (Hunyuan kong2024hunyuanvideo, Wan2.1-1.3B wan2025wan and OpenSora zheng2024open), and class-to-image generation (DiT-XL/2 with DDPM ho2020denoising and DDIM song2020denoising samplers). Our method is integrated into five SOTA acceleration methods: FasterSD li2023faster, TeaCache liu2025timestep, ToCa zou2024accelerating, DuCa zou2024accelerating1 and TaylorSeer liu2025reusing. In addition, we ensure compatibility with quantized models by applying our approach to the Q-DiT chen2025q.
@@
Evaluation and metrics. For text-to-image generation, we use captions of MS-COCO2017 lin2014microsoft to generate images and evaluate performance using FID heusel2017gans, CLIPScore (CLIP) hessel2021clipscore, ImageReward (IR) xu2023imagereward, Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS). Note that variations in sample counts (10K or 50K) and CLIP versions stem from differences in the respective acceleration baselines. For text-to-video generation, we report the average performance across 16 tasks from the VBench huang2024vbench. For the class-to-image generation, we randomly sample 50,000 class labels from ImageNet and compute FID, sFID, InceptionScore (IS) salimans2016improved, Precision (P) and Recall (R). Finally, we evaluate generation speed on RTX4090 or A800 using both FLOPs and latency. Refer to Appendix. D.1 for more details.
@@ 5 Conclusion
    A The Use of Large Language Models
@@ D More Experiments
    • D.4 Comparison with learning‑based methods
@@ E More Visualization
Appendix A The Use of Large Language Models
@@ C.1 Dynamic Programming Pseudocode
where $g$ models propagation (approximately linear due to DiT’s residual connections and attention linearity). The CUMSUM approximation assumes $g\approx\text{Id}$ (identity, i.e., direct summation), which holds because of “high structural similarity between input and output” (as noted in liu2025timestep).
@@ D.1 More Implementation Details
Generation models. Text-to-Image Generation: We evaluate three diffusion-based text-to-image generation models: (1) Stable Diffusion v1.5 (SD1.5) rombach2022high, a latent diffusion model (LDM) trained on LAION-5B, generating images at a resolution of $\times512$. (2) PixArt-$\bm{\alpha}$ chen2023pixart, a transformer-based diffusion model that performs efficient pixel-space modeling at $\times256$ resolution. (3) FLUX.1-dev flux2024, a state-of-the-art diffusion transformer trained at $\times1024$ resolution, capable of producing high-fidelity, photorealistic images.
@@
Text-to-Video Generation: We include three representative video diffusion models: (1) HunyuanVideo kong2024hunyuanvideo, generating 65 frames at 480p resolution, emphasizing realistic motion and temporal coherence. (2) Wan2.1-1.3B wan2025wan, a lightweight and efficient video diffusion transformer generating 65 frames at 480p. (3) OpenSora zheng2024open, an open research model producing 2-second clips at 480p, enabling temporally consistent text-conditioned generation.
@@
Class-to-Image Generation: We adopt DiT-XL/2 peebles2023scalable, a large diffusion transformer trained on ImageNet, evaluated with both DDPM ho2020denoising and DDIM song2020denoising samplers at a $\times256$ resolution.
@@
Acceleration Baselines. Our method (CEM) is integrated into six representative acceleration or efficiency improvement approaches: FasterSD li2023faster accelerates diffusion by reusing features extracted from shallow network layers to reduce redundant computation in deeper ones; ToCa zou2024accelerating combines caching and pruning mechanisms to jointly optimize computational reuse and step reduction; DuCa zou2024accelerating1 integrates conservative and aggressive caching strategies to maintain a balance between fidelity and acceleration; TaylorSeer liu2025reusing predicts reusable features from historical cache states through a Taylor-series expansion rather than directly reusing cached results; TeaCache liu2025timestep adaptively determines the caching policy based on the relationship between input and output activations; and Q-DiT chen2025q is a training-free quantized Diffusion Transformer that serves as our platform to validate the compatibility of CEM with quantized models.
@@
Evaluation and Metrics. For text-to-image generation, all models produce images conditioned on captions from the MS-COCO 2017 dataset lin2014microsoft. For text-to-video generation, evaluations are conducted on VBench huang2024vbench, a comprehensive benchmark of 16 sub-tasks assessing spatial fidelity, motion consistency, and text–video alignment. For class-to-image generation, DiT-XL/2 is evaluated on ImageNet deng2009imagenet by generating images for its 1,000 labeled categories.
@@
We employ standard metrics to assess fidelity, perceptual quality, diversity, and semantic alignment. Fréchet Inception Distance (FID) heusel2017gans measures distributional similarity between generated and real images; lower values indicate better fidelity. CLIPScore (CLIP) hessel2021clipscore evaluates semantic alignment between text and images, and higher scores denote stronger consistency. ImageReward (IR) xu2023imagereward reflects human aesthetic preference; higher is better. Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) evaluate pixel-level and structural fidelity, both maximized for better reconstruction quality. Learned Perceptual Image Patch Similarity (LPIPS) quantifies perceptual difference in deep feature space; lower indicates higher similarity.
@@
For video generation, VBench huang2024vbench jointly evaluates realism, temporal smoothness, and motion–text consistency, with higher scores representing better video quality. In class-to-image generation, we report Sliced FID (sFID) for image fidelity, Inception Score (IS) salimans2016improved for diversity, and Precision (P) and Recall (R) to measure the trade-off between fidelity and diversity.

### v2

# arXiv:2512.23258v2[cs.CV] 28 Feb 2026

@@ Abstract
Although Diffusion Transformer (DiT) has emerged as a predominant architecture for image and video generation, its iterative denoising process results in slow inference, which hinders broader applicability and development. Caching-based methods achieve training-free acceleration, while suffering from considerable computational error. Existing methods typically incorporate error correction strategies such as pruning or prediction to mitigate it. However, their fixed caching strategy fails to adapt to the complex error variations during denoising, which limits the full potential of error correction. To tackle this challenge, we propose a novel fidelity-optimization plugin for existing error correction methods via cumulative error minimization, named CEM. CEM predefines the error to characterize the sensitivity of model to acceleration jointly influenced by timesteps and cache intervals. Guided by this prior, we formulate a dynamic programming algorithm with cumulative error approximation for strategy optimization, which achieves the caching error minimization, resulting in a substantial improvement in generation fidelity. CEM is model-agnostic and exhibits strong generalization, which is adaptable to arbitrary acceleration budgets. It can be seamlessly integrated into existing error correction frameworks and quantized models without introducing any additional computational overhead. Extensive experiments conducted on nine generation models and quantized methods across three tasks demonstrate that CEM significantly improves generation fidelity of existing acceleration models, and outperforms the original generation performance on FLUX.1-dev, PixArt-$\alpha$, StableDiffusion1.5 and Hunyuan. Our code is released publicly at https://github.com/leaves162/CEM.
@@ 1 Introduction
Diffusion models Ho et al. (2020); Rombach et al. (2022) have significantly advanced visual generation tasks, including image Chen et al. (2023) and video Zheng et al. (2024) generation, and extended to other tasks Nie et al. (2025); Hu et al. (2025); Sun et al. (2025). Diffusion Transformers (DiT) Peebles and Xie (2023) have emerged as the dominant architecture for diffusion models, replacing U-Net Ronneberger et al. (2015) due to their inherent scalability and superior generative capabilities. Despite the impressive performance of these powerful models, their slow inference speed remains a critical barrier to widespread adoption. Currently, image generation typically requires tens of seconds, while video generation can take up to several minutes or even longer. This limitation primarily stems from the sequential nature of the reverse denoising process, which precludes parallel decoding. Moreover, the expansion of parameters, coupled with computationally intensive operations like attention Vaswani et al. (2017), further worsens efficiency.
To accelerate the DiT-based diffusion generation, researchers adopt techniques such as distillation Yin et al. (2024); Salimans and Ho (2022); Gu et al. (2025), learning Huang et al. (2024a); Ma et al. (2024a) and quantization Shang et al. (2023); He et al. (2023); Li et al. (2025), aiming to reduce inference latency by decreasing the model size. However, they necessitate training phases, which entails substantial computational cost, and lack generalization across models. An alternative approach is the caching mechanism Liu et al. (2025b); Saghatchian et al. (2025); Qiu et al. (2025), which leverages the similarity Ma et al. (2024b) between adjacent timesteps or layers to reuse previous cached hidden states, achieving training-free acceleration. Naive caching Selvaraju et al. (2024); Ma et al. (2024b) inevitably accumulates noise during the denoising, with error growing exponentially as the cache interval increases, resulting in a substantial deterioration in generation fidelity.
Therefore, existing caching optimization methods are combined with additional mechanisms to perform error correction, aiming to mitigate the cumulative caching error. For example, methods such as ToCa Zou et al. (2024a), DuCa Zou et al. (2024b), and FastCache Liu et al. (2025a) incorporate the pruning operation when reusing cache to retain the computation of a subset of important tokens, thereby reducing error. In addition, methods such as ICC Chen et al. (2025b) and TaylorSeer Liu et al. (2025c); Guan et al. (2025) leverage historical trends to predict the output variations across different timesteps and guide the update of cached representations, rather than directly reusing cache.
@@
Although the idea of error correction partially reduces the cumulative error for balancing the generation quality and acceleration efficiency, its effectiveness is limited by the error of relatively simple or fixed caching strategies. For example, ToCa Zou et al. (2024a) and DuCa Zou et al. (2024b) adopt the cache schedule that changes linearly with the timestep, while the cache interval of methods like ICC Chen et al. (2025b) and TaylorSeer Liu et al. (2025c); Guan et al. (2025) is a constant. They are incapable of handling the complex dynamics of sensitivity to caching during denoising. This limitation results in an insufficient characterization of the model’s intrinsic sensitivity to caching, thereby preventing the mitigation of error accumulation during acceleration. Consequently, the potential of error correction is limited, leading to degradation in generation fidelity.
@@
Although several prior works Qiu et al. (2025) like TeaCache Liu et al. (2025b) have explored caching optimization, they rely on real-time error estimation, introducing additional computational overhead that undermines efficiency gains. Furthermore, they are tied to specific acceleration ratios or model architectures, limiting compatibility with error correction. In contrast, our CEM conducts offline error modeling prior to inference, enabling the use of prior knowledge without incurring runtime cost. CEM is model-agnostic, supports arbitrary acceleration budgets, and integrates seamlessly with both error correction methods and quantized models.
@@
  • We introduce dynamic programming based on prior errors to derive the optimal caching strategy that minimizes cumulative error, thereby reducing the error of existing caching-based methods.
@@ 2 Related Work
Diffusion Transformer. Diffusion models Ho et al. (2020); Rombach et al. (2022) have demonstrated remarkable capabilities in image Chen et al. (2023) and video Zheng et al. (2024); Kong et al. (2024) generation. Upon this, DiT Peebles and Xie (2023) replace the convolutional modules in U-Net Ronneberger et al. (2015) with attention Vaswani et al. (2017), which enhances scalability and leads to breakthrough improvements in generation quality. Despite these advancements, the iterative nature of the diffusion combined with the computational complexity of attention results in slow inference. As a result, acceleration techniques have emerged as a key focus of research.
@@
DiT Acceleration. General acceleration techniques such as distillation Meng et al. (2023); Yao et al. (2024) and quantization Li et al. (2023b); Liu et al. (2025d) require retraining models, which incur substantial time costs and lacks generalization across models. Consequently, training-free methods have emerged as a promising direction. Caching-based methods Selvaraju et al. (2024); Zou et al. (2025); Lv et al. (2024) accelerate inference by reusing previous hidden states with high similarity, while the reuse of cache inevitably introduces error, which accumulate over time and result in a significant degradation of generation quality. To correct this error, two strategies have been proposed: (1) Token pruning, which retains the computation of a subset of tokens during cache reuse to reduce the error caused by fully relying on cache. For example, DuCa Zou et al. (2024b) applies pruning at specific timesteps to balance generation quality and acceleration efficiency. (2) Predictive reuse, which estimates future representations based on historical evolution of cache rather than directly reusing them. For instance, TaylorSeer Liu et al. (2025c); Guan et al. (2025) performs a taylor expansion on cache and utilizes gradient dynamics to predict features for reuse.
@@
These methods alleviate caching error to some extent, they overlook the optimization of the caching strategy itself, corrections are applied on top of relatively high cumulative error, limiting their effectiveness. Although a few studies have explored strategy optimization, they are difficult to integrate with error correction for quality enhancement. For instance, AdaCache Kahatapitiya et al. (2024) relies on motion trajectories and is restricted to video; AdaptiveDiffusion Ye et al. (2024) introduces overhead from third-order calculations and is specific to the U-Net; TeaCache Liu et al. (2025b) performs real-time estimation and polynomial fitting, while the extra cost offsets the benefits of cache-based acceleration. To address these issues, we propose a plug-and-play acceleration framework that optimizes the caching strategy by offline prior, seamlessly integrates with error correction methods without computational overhead, substantially improving their generation fidelity.
@@
In addition, optimizing the sampling strategy also leads to acceleration. For example, DDIM Song et al. (2020) introduces deterministic sampling to reduce the number of denoising steps. DPM-Solver Lu et al. (2022) and DPM-Solver++ Lu et al. (2025) leverage adaptive high-order solvers to accelerate denoising. Rectified flow Liu et al. (2022); Liu (2022) improves the transport of the distribution in the ordinary differential equation. Our method is orthogonal to these approaches and can be integrated across sampling strategies, even supporting compatibility with quantized models.
@@ 3.1 Offline Error Modeling
Joint modeling analysis. We model the error as a joint function of denoising timesteps and cache intervals. Previous methods Zou et al. (2024a; b); Liu et al. (2025c) only consider differences across timesteps and determine cache intervals heuristically, limiting control over the acceleration budget. In contrast, we explicitly model cache intervals together with timesteps to capture the error distribution and analyze the sensitivity of different denoising stages. This formulation is simple yet accurate, offering flexible acceleration control, while its offline design introduces no additional computational overhead and provides the foundation for our algorithm of cumulative error minimization in Sec. 3.2.
@@ 4.1 Experiment Settings
Models and baselines. We conduct extensive experiments across three tasks: text-to-image generation (using StableDiffusion1.5 (SD1.5) Rombach et al. (2022), PixArt-$\bm{\alpha}$ Chen et al. (2023) and FLUX.1-dev Labs (2024)), text-to-video generation (Hunyuan Kong et al. (2024), Wan2.1-1.3B Wan et al. (2025) and OpenSora Zheng et al. (2024)), and class-to-image generation (DiT-XL/2 with DDPM Ho et al. (2020) and DDIM Song et al. (2020) samplers). Our method is integrated into five SOTA acceleration methods: FasterSD Li et al. (2023a), TeaCache Liu et al. (2025b), ToCa Zou et al. (2024a), DuCa Zou et al. (2024b) and TaylorSeer Liu et al. (2025c). In addition, we ensure compatibility with quantized models by applying our approach to the Q-DiT Chen et al. (2025a).
@@
Evaluation and metrics. For text-to-image generation, we use captions of MS-COCO2017 Lin et al. (2014) to generate images and evaluate performance using FID Heusel et al. (2017), CLIPScore (CLIP) Hessel et al. (2021), ImageReward (IR) Xu et al. (2023), Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS). Note that variations in sample counts (10K or 50K) and CLIP versions stem from differences in the respective acceleration baselines. For text-to-video generation, we report the average performance across 16 tasks from the VBench Huang et al. (2024b). For the class-to-image generation, we randomly sample 50,000 class labels from ImageNet and compute FID, sFID, InceptionScore (IS) Salimans et al. (2016), Precision (P) and Recall (R). Finally, we evaluate generation speed on RTX4090 or A800 using both FLOPs and latency. Refer to Appendix. D.1 for more details.
@@ 5 Conclusion
Acknowledgments
This work was supported by National Natural Science Foundation of China (grant No. 62350710797), by Guangdong Basic and Applied Basic Research Foundation (grant No. 2023B1515120065, 2025A1515011546), and by the Shenzhen Science and Technology Program (JCYJ20240813105901003, KJZD20240903102901003, ZDCY20250901113000001).
@@ Acknowledgments
    A Statements
    • A.1 Ethics Statement
    • A.2 Reproducibility Statement
    • A.3 The Use of Large Language Models
@@ D More Experiments
    • D.4 Comparison with Learning-based Methods
@@ E More Visualization
Appendix A Statements
A.1 Ethics Statement
We affirm that the research and writing of this paper strictly adhere to the ICLR Code of Ethics, and do not involve any issues related to human subjects, ethical concerns, or potential conflicts of interest.
A.2 Reproducibility Statement
We affirm that all research presented in this paper is fully reproducible, including the error modeling analysis in Sec. 3.1, the core implementation described in Sec. 3.2 and Appendix LABEL:sec:dp-details, and all results and visualizations shown in figures and tables. The source code will be released publicly later.
A.3 The Use of Large Language Models
@@ C.1 Dynamic Programming Pseudocode
where $g$ models propagation (approximately linear due to DiT’s residual connections and attention linearity). The CUMSUM approximation assumes $g\approx\text{Id}$ (identity, i.e., direct summation), which holds because of “high structural similarity between input and output” (as noted in Liu et al. (2025b)).
@@ D.1 More Implementation Details
Generation models. Text-to-Image Generation: We evaluate three diffusion-based text-to-image generation models: (1) Stable Diffusion v1.5 (SD1.5) Rombach et al. (2022), a latent diffusion model (LDM) trained on LAION-5B, generating images at a resolution of $\times512$. (2) PixArt-$\bm{\alpha}$ Chen et al. (2023), a transformer-based diffusion model that performs efficient pixel-space modeling at $\times256$ resolution. (3) FLUX.1-dev Labs (2024), a state-of-the-art diffusion transformer trained at $\times1024$ resolution, capable of producing high-fidelity, photorealistic images.
@@
Text-to-Video Generation: We include three representative video diffusion models: (1) HunyuanVideo Kong et al. (2024), generating 65 frames at 480p resolution, emphasizing realistic motion and temporal coherence. (2) Wan2.1-1.3B Wan et al. (2025), a lightweight and efficient video diffusion transformer generating 65 frames at 480p. (3) OpenSora Zheng et al. (2024), an open research model producing 2-second clips at 480p, enabling temporally consistent text-conditioned generation.
@@
Class-to-Image Generation: We adopt DiT-XL/2 Peebles and Xie (2023), a large diffusion transformer trained on ImageNet, evaluated with both DDPM Ho et al. (2020) and DDIM Song et al. (2020) samplers at a $\times256$ resolution.
@@
Acceleration Baselines. Our method (CEM) is integrated into six representative acceleration or efficiency improvement approaches: FasterSD Li et al. (2023a) accelerates diffusion by reusing features extracted from shallow network layers to reduce redundant computation in deeper ones; ToCa Zou et al. (2024a) combines caching and pruning mechanisms to jointly optimize computational reuse and step reduction; DuCa Zou et al. (2024b) integrates conservative and aggressive caching strategies to maintain a balance between fidelity and acceleration; TaylorSeer Liu et al. (2025c) predicts reusable features from historical cache states through a Taylor-series expansion rather than directly reusing cached results; TeaCache Liu et al. (2025b) adaptively determines the caching policy based on the relationship between input and output activations; and Q-DiT Chen et al. (2025a) is a training-free quantized Diffusion Transformer that serves as our platform to validate the compatibility of CEM with quantized models.
@@
Evaluation and Metrics. For text-to-image generation, all models produce images conditioned on captions from the MS-COCO 2017 dataset Lin et al. (2014). For text-to-video generation, evaluations are conducted on VBench Huang et al. (2024b), a comprehensive benchmark of 16 sub-tasks assessing spatial fidelity, motion consistency, and text–video alignment. For class-to-image generation, DiT-XL/2 is evaluated on ImageNet Deng et al. (2009) by generating images for its 1,000 labeled categories.
@@
We employ standard metrics to assess fidelity, perceptual quality, diversity, and semantic alignment. Fréchet Inception Distance (FID) Heusel et al. (2017) measures distributional similarity between generated and real images; lower values indicate better fidelity. CLIPScore (CLIP) Hessel et al. (2021) evaluates semantic alignment between text and images, and higher scores denote stronger consistency. ImageReward (IR) Xu et al. (2023) reflects human aesthetic preference; higher is better. Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) evaluate pixel-level and structural fidelity, both maximized for better reconstruction quality. Learned Perceptual Image Patch Similarity (LPIPS) quantifies perceptual difference in deep feature space; lower indicates higher similarity.
@@
For video generation, VBench Huang et al. (2024b) jointly evaluates realism, temporal smoothness, and motion–text consistency, with higher scores representing better video quality. In class-to-image generation, we report Sliced FID (sFID) for image fidelity, Inception Score (IS) Salimans et al. (2016) for diversity, and Precision (P) and Recall (R) to measure the trade-off between fidelity and diversity.